### v1

@@ Abstract
The practical channel estimation in uplink pinching-antenna systems is investigated, in which an electromagnetic-compliant in-waveguide transmission model is exhibited, incorporating both bidirectional power splitting, cumulative power leakage, and waveguide attenuation. Based on this model, the paper investigates two antenna activation protocols for channel estimation: a serial protocol based on one-by-one antenna activation and a parallel protocol utilizing a binary S-Matrix activation. The serial protocol is characterized by its superior numerical stability but a lack of array gain, whereas the parallel protocol theoretically offers array gain but suffers from severe performance degradation due to structural crosstalk from the non-orthogonal S-Matrix and ill-conditioning from cumulative leakage. Furthermore, the paper analyzes the fundamental commonalities and asymmetries between uplink and downlink channel estimation in pinching-antenna systems. Numerical results demonstrate that 1) in an ideal lossless model, the parallel protocol is superior to the serial protocol due to the array gain from simultaneous energy collection in uplink transmission; 2) in a practical model with physical losses, the serial protocol outperforms the parallel protocol, as the performance of the parallel protocol is degraded by the numerical instability from cumulative leakage, which outweighs the benefit of array gain; 3) For downlink channel estimation, the serial protocol is more suitable because it avoids bidirectional power splitting, while the parallel protocol is more suitable for the uplink as it can make full use of array gain.
Index Terms:
Pinching antenna, channel estimation
@@ II System Models
For clarity of derivation and without loss of generality, we first analyze the case of a single waveguide. The results can be directly extended to the multi-waveguide case as the estimation on each waveguide can be considered an independent process. We consider a AP is equipped with a waveguide to serve $K$ single-antenna user equipments (UEs). The waveguide contains $N$ controllable PAs. The waveguide are installed at a height $d$ and are parallel to the $x$-axis. In a three-dimensional Cartesian system, the UEs are assumed to be randomly distributed within a rectangular region on the $x$-$y$ plane, with dimensions $D_{x}$ and $D_{y}$. The position of the $k$-th UE is represented by $\psi_{k}=(x_{k},y_{k},0)$.
@@ III-A Problem Formulation
The ultimate goal of uplink channel estimation is to recover the wireless channel vector $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, which captures the environment-dependent propagation characteristics and is time-varying. In contrast, the in-waveguide transfer vectors $\mathbf{g}_{\text{serial}}$ and $\mathbf{g}_{\text{parallel}}$ are hardware-dependent, determined by waveguide geometry, PA spacing, and material properties, and can be pre-calibrated offline, thus treated as known prior to estimation. We assume that time-orthogonal pilot sequences are assigned to the $K$ users, which allows for interference-free channel estimation for each user individually. The subsequent analysis focuses on the estimation process for an arbitrary user $k$. Therefore, for notational simplicity, the user index $k$ is dropped from all relevant terms. The derived performance metrics are applicable to any user in the system. We define the composite channel $\mathbf{b}$ for each operational mode. For serial estimation, we define
@@ IV-A1 MSE of Serial Estimator for $\mathbf{h}$
Since $\mathbf{G}_{\text{serial}}$ is diagonal, ${\mathbf{G}_{\text{serial}}}^{H}\mathbf{G}_{\text{serial}}$ is also diagonal with elements $|[g_{\text{serial}}]_{n}|^{2}=\gamma e^{-2\alpha d_{n}}$. The inverse of a diagonal matrix is diagonal with elements equal to the inverse of the original diagonal elements, we have
@@ VI Numerical Results
Fig. 3 provides deeper insight into the uplink performance trade-off by presenting the NMSE as a function of the pass-through coefficient $\beta$. The performance of the practical serial protocol appears as a horizontal line, confirming its independence from inter-PA leakage and establishing it as a robust baseline. In stark contrast, the proportional radiation model shows extreme sensitivity to $\beta$. As $\beta$ decreases, the observation matrix becomes rapidly ill-conditioned, leading to catastrophic noise amplification and an exponential degradation in NMSE. The equal radiation model with numerical stability maintains the stale performance across nearly the entire range of $\beta$, while its performance slightly degrades as $\beta\to 1$ because the corresponding coupling coefficient $\alpha=\sqrt{1-\beta^{2}}$ approaches zero, reducing the total energy coupled into the system.
@@
Fig. 5 shows the NMSE performance of for the downlink channel estimation. The downlink estimation is fundamentally a power-budget-limited problem, where the serial protocol can concentrate the entire transmit power of the access point on a single PA measurement, thereby maximizing its SNR. Compared to uplink channel estimation, the downlink channel estimation can also avoid the bidirectional power splitting. In contrast, the parallel protocols must divide the total power among $N$ PAs, incurring an irrecoverable $1/N$ power allocation loss. Consequently, the serial protocol is the clear winner among all schemes in the downlink channel estimation. Note that the channel feedback overhead needs to be considered if the channel estimation is carried out at the user side.

### v2

@@ Abstract
The practical channel estimation for pinching-antenna networks is investigated, in which an electromagnetic-compliant in-waveguide transmission model is exhibited, incorporating bidirectional power splitting, cumulative power leakage, and waveguide attenuation. Based on this model, the paper investigates two antenna activation protocols for channel estimation: a serial protocol based on one-by-one antenna activation and a parallel protocol utilizing a binary S-Matrix activation. The serial protocol is characterized by its superior numerical stability but a lack of array gain, whereas the parallel protocol theoretically offers array gain but suffers from severe performance degradation due to structural crosstalk from the non-orthogonal S-Matrix and ill-conditioning from cumulative leakage. Furthermore, the paper analyzes the fundamental commonalities and asymmetries between uplink and downlink channel estimation in pinching-antenna systems. Numerical results demonstrate that 1) in an ideal lossless model, the parallel protocol is superior to the serial protocol due to the array gain from simultaneous energy collection in uplink transmission; 2) in a practical model with physical losses, the serial protocol outperforms the parallel protocol, as the performance of the parallel protocol is degraded by the numerical instability from cumulative leakage, which outweighs the benefit of array gain; 3) For downlink channel estimation, the serial protocol is more suitable because its strategy of concentrating the entire power budget on one measurement, while the parallel protocol is more suitable for the uplink as it can make full use of array gain.
@@ II System Models
For clarity of derivation and without loss of generality, we first analyze the case of a single waveguide. The results can be directly extended to the multi-waveguide case as the estimation on each waveguide can be considered an independent process. We consider an AP is equipped with a waveguide to serve $K$ single-antenna user equipments (UEs). The waveguide contains $N$ controllable PAs. The waveguide is installed at a height $d$. In a three-dimensional Cartesian system, the UEs are assumed to be randomly distributed within a rectangular region on the $x$-$y$ plane, with dimensions $D_{x}$ and $D_{y}$. The position of the $k$-th UE is represented by $\psi_{k}=(x_{k},y_{k},0)$.
@@ III-A Problem Formulation
The ultimate goal of uplink channel estimation is to recover the wireless channel vector $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$, which captures the environment-dependent propagation characteristics and is time-varying. In contrast, the in-waveguide transfer vectors $\mathbf{g}_{\text{serial}}$ and $\mathbf{g}_{\text{parallel}}$ are hardware-dependent, determined by waveguide geometry, PA spacing, and material properties, and can be pre-calibrated offline, thus treated as known prior to estimation. We assume that time-orthogonal pilot sequences are assigned to the $K$ users, which allows for interference-free channel estimation for each user individually. In this work, the subsequent analysis focuses on the estimation process for an arbitrary user $k$ and the initial channel estimation for the currently activated antennas. Therefore, for notational simplicity, the user index $k$ is dropped from all relevant terms. The derived performance metrics are applicable to any user in the system. We define the composite channel $\mathbf{b}$ for each operational mode. For serial estimation, we define
@@ IV-A1 MSE of Serial Estimator for $\mathbf{h}$
Since $\mathbf{G}_{\text{serial}}$ is diagonal, ${\mathbf{G}_{\text{serial}}}^{H}\mathbf{G}_{\text{serial}}$ is also diagonal with elements $|[g_{\text{serial}}]_{n}|^{2}=\gamma e^{-2\alpha d_{n}}$. The inverse of a diagonal matrix is diagonal with elements equal to the inverse of the original diagonal elements. Thus, we have
@@ VI Numerical Results
Fig. 3 provides deeper insight into the uplink performance trade-off by presenting the NMSE as a function of the pass-through coefficient $\beta$. The performance of the practical serial protocol appears as a horizontal line, confirming its independence from inter-PA leakage and establishing it as a robust baseline. In stark contrast, the proportional radiation model shows extreme sensitivity to $\beta$. As $\beta$ decreases, the observation matrix becomes rapidly ill-conditioned, leading to catastrophic noise amplification and an exponential degradation in NMSE. The equal radiation model with numerical stability maintains the stable performance across nearly the entire range of $\beta$, while its performance slightly degrades as $\beta\to 1$ because the corresponding coupling coefficient $\alpha=\sqrt{1-\beta^{2}}$ approaches zero, reducing the total energy coupled into the system.
@@
Fig. 5 shows the NMSE performance for the downlink channel estimation. The downlink estimation is fundamentally a power-budget-limited problem, where the serial protocol can concentrate the entire transmit power of the access point on a single PA measurement, thereby maximizing its SNR. Compared to uplink channel estimation, the downlink channel estimation can also avoid the bidirectional power splitting. In contrast, the parallel protocols must divide the total power among $N$ PAs, incurring an irrecoverable $1/N$ power allocation loss. Consequently, the serial protocol is the clear winner among all schemes in the downlink channel estimation. Note that the channel feedback overhead needs to be considered if the channel estimation is carried out at the user side.